# Gravitational quasinormal modes of black holes in quadratic gravity

### v1

@@ I Introduction
Arguably, the most natural way to modify GR is by introducing higher derivative operators to the Einstein-Hilbert action Berti et al. (2015). Such terms may be motivated either from an effective field (EFT) theory point of view or from more fundamental arguments. For example, in so-called quadratic gravity the gravitational action is modified by including terms that are quadratic in the curvature tensors. Such terms make the theory formally renormablizable Stelle (1977) but introduce ghosts, associated with extra degrees of freedom of massive spin-0 and spin-2 modes propagating in this theory. Thus, such theory can be only considered viable as an EFT, and only for physical processes at energy lower than the cut-off scale. Remarkably, despite the presence of ghosts and the fourth-order field equations stemming from its action, quadratic gravity was found to lead to a well-posed initial-value formulation Noakes (1983). Certain subclasses of quadratic gravity have gathered particular interest in the context of inflationary cosmology (e.g., Starobinski’s model of inflation Starobinsky (1980)) since quadratic curvature corrections might be relevant in the early universe.
Another natural arena to explore the implications of quadratic curvature terms are black holes (BHs) Barack et al. (2019). As long as we consider perturbations of static, asymptotically flat spacetimes, it can be shown Nelson (2010); Lu et al. (2015) that it is not restrictive to consides a subclass of quadratic gravity theories, the so-called Einstein-Weyl (EW) gravity, in which the Einstein-Hilbert term is supplemented by the square of the Weyl tensor. This term has a dimensionful coupling constant $\alpha$, which is associated with the mass $\mu$ of the massive spin-2 degree of freedom propagating in this theory, namely $\alpha\equiv 1/(2\mu^{2})$ (hereafter we use $G=c=1$ units).
@@
The horizon radius $r_{h}$ of the hairy BH is bounded both from above and from below. BH solutions with $r_{h}$ larger than the upper bound have negative mass and are thus unphysical. Indeed, at variance with the Schwarzschild solution, the mass of the hairy BH is a decreasing function of the horizon radius. BH solutions with $r_{h}$ smaller than the lower bound are unstable against radial perturbations Held and Zhang (2023). In terms of the dimensionless parameter $p\equiv r_{h}/\sqrt{2\alpha}$, radially stable hairy BHs with a positive mass exist for $0.876\lesssim p\lesssim 1.143$. In the case of stationary, rotating BHs, an approximate computation shows the same structure of the solution space: the Kerr solution is stable above a critical value of mass, while a hairy BH solution exists only below this value Sajadi and Hendi (2023).
In this paper we consider non-radial perturbations of both branches of spherically symmetric BH solutions in EW gravity, studying its quasi-normal modes (QNMs) of oscillations, which describe the damped sinusoids corresponding to the free oscillations of a perturbed BH.
@@
Secondly, the QNM spectrum is a promising probe to observationally test if GR deviations of the class of those studied in this paper are actually present. QNMs play a crucial role in the post-merger ringdown phase at the end of a binary BH coalescence and encapsulate important information about the underlying gravitational theory (see Berti et al. (2018) for a review). Today, QNMs have been measured for a number of binary BH coalescences Ghosh et al. (2021); Abbott et al. (2021b). In GR, the QNMs have been extensively studied and have been shown to be determined by the mass, charge, and spin of the BH, in accordance with no-hair theorems. However, in theories beyond GR, the ringdown spectrum exhibits distinct features that can be used to distinguish them from GR. In particular, any modified gravity theory generically introduces two modifications to the ringdown: i) a deformed QNM spectrum, whose deviations from the GR case are proportional to (powers of) the coupling constant; ii) novel extra modes in the gravitational waveform D’Addario et al. (2024); Crescimbeni et al. (2024), whose amplitude is proportional to (powers of) the coupling.
@@
BHs for $p>0.876$, instead, are stable for radial perturbations. Time-domain1 evolutions of BHs in the Schwarzschild branch Held and Lim (2021), give strong indications for their non-radial stability (see also Held and Lim (2023) for the evolution of rotating BHs and of BH binaries). Moreover, scalar and electromagnetic perturbations of BHs in the non-GR branch have been studied using approximate techniques, finding stable modes Cai et al. (2016); Zinhailo (2018, 2019); Konoplya (2023). In this paper we compute for the first time gravitational QNMs, in the axial sector, for both branches of spherically symmetric BH solutions in EW gravity, finding as expected a deformation of the GR spectrum and new classes of modes.
Finally we study whether the new modes can be excited in the ringdown stage of a binary BH coalescence. Then, using the results of our study, we discuss the phenomenological perspectives of EW gravity, i.e. whether we expect the non-GR features of BHs in this theory, if present, could be observed by present or near-future GW detectors.
The paper’s structure is the following: in Sec. II we present the theoretical framework of the model. In Sec. III we review the background solutions in this theory, both with the numerical and semi-analytical treatment. In Sec. IV we analyze the axial perturbations in EW gravity, first on a Schwarzschild background recovering the results presented in Brito et al. (2013). We show our analysis for the axial QNMs of the non-GR branch of solutions in IV.2. In Sec. V we discuss the phenomenological perspectives of EW gravity. Finally, we draw our conclusions in Sec. VI.
@@ II Framework
where the GB invariant is defined as $\mathscr{G}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. The dimensionful parameters $\alpha$ and $\beta$ are associated with additional massive degrees of freedom. Specifically, in addition to the massless spin-2 graviton, this theory predicts a massive spin-0 mode with mass $m^{2}=1/(6\beta)$, and a massive spin-2 mode with mass $\mu^{2}=1/(2\alpha)$. Remarkably, the sign of the kinetic term of the massive spin-2 mode is the opposite than those of the other kinetic terms, i.e. it is a ghost-like degree of freedom. This makes quadratic gravity non-unitary as a full quantum theory: it can only be considered as a low-energy EFT, as long as the energy scale is low enough. We remark, however, that the existence of an Ostrogradsky ghost does not prevent the theory from having a well-posed initial value formulation at a classical level Noakes (1983); Held and Lim (2021).
@@
where beyond GR contributions are proportional to a factor $\alpha$. Despite not being obvious at first glance, the field equations for spherically symmetric configurations yield second order equations. We will go over that in the following subsections.
@@ II.1 Auxiliary field
$\begin{split}&\mathcal{E}^{(f)}_{\mu\nu}\equiv G_{\mu\nu}+G_{(\mu}{}^{\rho}f_{\nu)\rho}-g_{\mu\nu}G^{\rho\sigma}f_{\rho\sigma}-fR_{\mu\nu}\\ &+Rf_{\mu\nu}+\Box f_{\mu\nu}+(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu})f-2\nabla_{\rho}\nabla_{(\mu}f_{\nu)}{}^{\rho}\\ &+g_{\mu\nu}\nabla_{\rho}\nabla_{\sigma}f^{\rho\sigma}+\mu^{2}\big{[}\big{(}f^{2}-f_{\rho\sigma}f^{\rho\sigma}\big{)}/2\\ &+2\big{(}f_{\mu}{}^{\rho}f_{\nu\rho}-f\,f_{\mu\nu}\big{)}\big{]}=0\end{split}$ (10)
We remark that the GR limit corresponds to $\mu\to\infty$, i.e. $\alpha\to 0$. In this limit, the auxiliary field $f_{\mu\nu}$ (Eq. (8)) vanishes, and Eqs.(10) reduce to Einstein’s equations in vacuum. Conversely, in the $\mu\to 0$ limit the theory becomes strongly coupled, the action only have quadratic curvature terms, and EW gravity reduces to the pure Weyl gravity Weyl (1918).
@@ III Background solutions
From Eq. (9), using the symmetry of $f_{\mu\nu}$, we find $\nabla_{\mu}\nabla_{\nu}f^{\mu\nu}=\nabla_{\mu}\nabla_{\nu}f^{\nu\mu}=\Box f$. From the trace of (10) we find then $R=0$: the spherically symmetric solutions (in vacuum) of EW gravity are necessarily Ricci-scalar flat. This also means that from (7) the trace of the auxiliary field vanishes, i.e. $f=0$. From this property we derive the first equation for the metric functions, i.e.
@@
Given that the static and spherically symmetric solutions have necessarily $R=0=f$, we will divide the analysis into two types of BH solutions:
@@
Obviously the first type of solution is analytical, while the other must be found numerically solving the system of ordinary differential equations with suitable boundary conditions.
@@ III.1 Expansions at the horizon
To find the coefficients in the expansions we substitute them in the background equations , which we then solve order by order. We show here the first few coefficients
@@ III.3 Semi-analytical background
In Fig. 1 we present, as an example, the $(tt)$ and $(rr)$ metric components as functions of $r/r_{h}$, for the hairy BH solution with $p=1.1$. We show both the solution computed solving numerically the field equations, and the semi-analytical approximation up to fourth order in the continued fractions. For the numerical solution the process we follow is a shooting approach with respect to the non-GR parameter $b_{1}$, integrating from the horizon outwards, so that at asymptotic infinity the exponentially diverging solution vanishes. For comparison, we also show the Schwarzschild solution.
@@
III.4 domain1 of existence
The second branch of non-GR BH solutions can be found for Kokkotas et al. (2017)
@@
where we remind the reader that the dimensionless parameter $p$ is defined as $p\equiv{r_{h}}/{\sqrt{2\alpha}}\equiv r_{h}\mu$ In Fig. 2 we show the domain1 of existence and of radial stability of hairy BHs in EW gravity, corresponding to the decreasing curve. The vertical dashed line highlights the allowed range, which starts at the bifurcation point from the Schwarzschild branch, and ends at the point where solutions are characterized by a zero mass. The straight line, instead, correspond of the solutions of the Schwarzschild branch. The axes correspond to dimensionless quantities.
@@
Figure 2: domain1 of existence of BHs in EW gravity. The bifurcation point from the GR solution is found at $p_{\text{min}}\simeq 0.876$, while the the upper value of the positive mass range is located at $p_{\text{max}}\simeq 0.876$. The dashed part of the curves corresponds to solutions either radially unstable, or having negative mass.
@@ IV Perturbations
where $\varepsilon$ is a bookkeeping parameter, $\bar{g}_{\mu\nu}$ and $\bar{f}_{\mu\nu}$ are the background metric and auxilliary field corresponding to a static, spherically symmetric BH solution, and $\delta g_{\mu\nu}$, $\delta f_{\mu\nu}$ are their respective perturbations. Hereafter, any quantity evaluated on the non-perturbed background will be denoted with a bar. At first order in the perturbations, Eq. (6) gives:
@@ IV.1 Schwarzschild background
From the above, it is evident that the massive perturbations $\delta f_{\mu\nu}$ decouple from $\delta g_{\mu\nu}$. Indeed, setting $\delta f_{\mu\nu}=0$, Eq. (34) is satisfied, and Eq. (33) reduces to the standard Regge-Wheeler equation Regge and Wheeler (1957), leading to the QNMs of Schwarzschild BHs in GR. When $\delta f_{\mu\nu}\neq 0$, Eqs. (33)-(34) yield a set of equations for the massive spin-two degrees of freedom, which coincide with those presented in Brito et al. (2013) after substituting the Schwarzschild metric. Although those perturbations are massive and do propagate at infinity, they nevertheless source ordinary metric perturbations through Eq. (33). Axial perturbations of the massive spin-2 excitation $\delta f_{\mu\nu}$ can be decomposed in scalar tensor harmonics (see Appendix B), leading - for each value of the harmonic indexes - to two vector perturbations, $F_{0}(r)$, $F_{1}(r)$, and one tensor perturbation, $F_{2}(r)$. In the massless limit, the only dynamical perturbation in the decomposition of $\delta f_{\mu\nu}$ is the tensor perturbation $F_{2}$, while the vector ones become pure gauge  [^1] Note that in this limit one does not recover, for $\delta f_{\mu\nu}$, the same decomposition as the (axial) massless field $\delta g_{\mu\nu}$, because the tensorial decomposition of $\delta f_{\mu\nu}$ is different from the Regge-Wheeler gauge employed in the decomposition of $\delta g_{\mu\nu}$. In particular, on a Schwarzschild background $\delta f_{\mu\nu}$ is traceless and transverse. .
@@
$\big{[}r\left(r\omega^{2}-BA^{\prime}\right)-A\left(rB^{\prime}-2B+\mu^{2}r^{2}+\Lambda-2\right)\big{]}=0\,,$ (38)
where $\Lambda\equiv\ell(\ell+1)$. Finally, from Eq. (33), by eliminating $h_{0}$, we deduce
@@
We may now eliminate $F_{0}$ by making use of (35), in which case the decoupled equations for the massive perturbations can be written as a system of Schroödinger-type equations
@@
The terms $-A\mu^{2}$ in $\hat{V}_{11}$ and $\hat{V}_{22}$ are an indication that $Q$ and $Z$ correspond to massive modes propagating with mass $\mu^{2}$.
Notice that with these redefinitions for $F_{1}$ and $F_{2}$ we cannot simultaneously describe all three perturbation functions $(F_{1},F_{2},h_{1})$ in a Schrödinger-type system due to the presence of the $F_{2}^{\prime}$ term in (39). In order to study the full QNM spectrum, we solved the following system describing both massless and massive degrees of freedom:
@@
For ingoing waves at the horizon, ${\Psi}^{(j)}(r)\sim e^{-i\omega r_{*}}$. Therefore we expand
$h_{1}(r)=$ $\;e^{-i\omega r}(r-2M)^{-2iM\omega}\sum_{n=0}\tilde{h}_{1}^{(n)}(r-2M)^{n-1}\,.$ (47)
@@
Summarizing, for each value of the BH mass (i.e. of the horizon radius), for each $\ell$, we find three series of modes with axial parity, parametrized, as customarily by the principal number $n$: (i) the massless modes, which coincide with those of GR; (ii) the massive vector modes; (ii) the massive tensor modes.
We present our results in Figs. 3 4, where we show the fundamental ($n=0$) $\ell=2$ massive vector and tensor modes, respectively, computed using the DI approach with $n_{\text{max}}=4,8$, and the CF approach. We find that the DI results, increasing $n_{max}$, converge towards the CF results, which we use as a benchmark. The vector mode computed using the DI approach with $n_{\text{max}}=4$ is a very good approximation, within $1\%$, of the CF result for $p\lesssim 0.6$, while for larger values of $p$ we need $n_{\max}=8$ to reach the same level of accuracy. For the tensor mode, the DI approach with $n_{\text{max}}=4$ is accurate within $1\%$ for $p\lesssim 0.6$, while that with $n_{\text{max}}=8$ has the same accuracy for $p\lesssim 0.75$. In Fig. 4 we also show the fundamental massless mode, which also belongs to the spectrum, and is the same as in GR. It coincides with the $\mu\to 0$ limit of the massive tensor mode, while we remind that the massive vector mode is not dynamical in this limit (see Appendix B).

### v3

@@ I Introduction
Arguably, the most natural way to modify GR is by introducing higher derivative operators to the Einstein-Hilbert action Berti et al. (2015). Such terms may be motivated either from an effective field (EFT) theory point of view or from more fundamental arguments. For example, in so-called quadratic gravity the gravitational action is modified by including terms that are quadratic in the curvature tensors. Such terms make the theory formally renormablizable Stelle (1977) but introduce ghosts, associated with extra degrees of freedom of massive spin-0 and spin-2 modes propagating in this theory.
These ghosts may challenge the viability of quadratic gravity as a fundamental theory, even though this is presently matter of debate (see e.g. Donoghue and Menezes (2022)). On the other hand, it can be shown (see e.g. Burgess (2004); Endlich et al. (2017)) that by considering quadratic gravity as an EFT, the deviations from GR can be transformed by a field redefinition to higher-order terms.
Despite the presence of ghosts and the fourth-order field equations stemming from its action, quadratic gravity was found to lead to a well-posed initial-value formulation Noakes (1983). This has motivated recent phenomenological and numerical studies on this theory beyond the EFT approach, at least at the classical level Lu et al. (2015); Lü et al. (2015); Kokkotas et al. (2017); Konoplya (2023); East and Siemonsen (2023); Held and Lim (2023). Furthermore, certain subclasses of quadratic gravity have gathered particular interest in the context of inflationary cosmology (e.g., Starobinski’s model of inflation Starobinsky (1980)) since quadratic curvature corrections might be relevant in the early universe.
In this paper we will take the same perspective of recent studies Lu et al. (2015); Lü et al. (2015); Kokkotas et al. (2017); Konoplya (2023); East and Siemonsen (2023); Held and Lim (2021, 2023), and consider quadratic gravity as a full theory. When needed, we will highlight the differences and the common regimes of validity with the EFT. As we are going to discuss, some of our results are also relevant for an analysis in an EFT framework, possibly at the same order as quartic terms in the curvature tensor such as those studied in Endlich et al. (2017); Cardoso et al. (2018).
A natural arena to explore the implications of quadratic curvature terms are black holes (BHs) Barack et al. (2019). As long as we consider perturbations of static, asymptotically flat spacetimes, it can be shown Nelson (2010); Lu et al. (2015) that it is not restrictive to consider a subclass of quadratic gravity theories, the so-called Einstein-Weyl (EW) gravity, in which the Einstein-Hilbert term is supplemented by the square of the Weyl tensor. This term has a dimensionful coupling constant $\alpha$, which is associated with the mass $\mu$ of the massive spin-2 degree of freedom propagating in this theory, namely $\alpha\equiv 1/(2\mu^{2})$ (hereafter we use $G=c=1$ units). The GR limit wherein an EFT approach is valid corresponds to $\alpha\to 0$ or, equivalently, $\mu\to\infty$, where the extra modes becomes very heavy.
@@
The horizon radius $r_{h}$ of the hairy BH is bounded both from above and from below. BH solutions with $r_{h}$ larger than the upper bound have negative mass and are thus unphysical. Indeed, at variance with the Schwarzschild solution, the mass of the hairy BH is a decreasing function of the horizon radius. BH solutions with $r_{h}$ smaller than the lower bound are unstable against radial perturbations Held and Zhang (2023). In terms of the dimensionless parameter $p\equiv r_{h}/\sqrt{2\alpha}$, radially stable hairy BHs with a positive mass exist for $0.876\lesssim p\lesssim 1.143$. The upper bound on $p$ in the domain of existence suggests that hairy BHs are non-perturbative solutions which do not exist within the EFT.
In the case of stationary, rotating BHs, an approximate computation shows the same structure of the solution space: the Kerr solution is stable above a critical value of mass, while a hairy BH solution exists only below this value Sajadi and Hendi (2023).
In this paper we consider non-radial perturbations of both branches of spherically symmetric BH solutions in EW gravity, studying its quasi-normal modes (QNMs) of oscillation, which describe the damped sinusoids corresponding to the free oscillations of a perturbed BH.
@@
Secondly, the QNM spectrum is a promising probe to observationally test if GR deviations of the class of those studied in this paper are actually present. QNMs play a crucial role in the post-merger ringdown phase at the end of a binary BH coalescence and encapsulate important information about the underlying gravitational theory (see Berti et al. (2018) for a review). Today, QNMs have been measured for a number of binary BH coalescences Ghosh et al. (2021); Abbott et al. (2021b). In GR, the QNMs have been extensively studied and have been shown to be determined by the mass, charge, and spin of the BH, in accordance with no-hair theorems. However, in theories beyond GR the ringdown spectrum exhibits distinct features that can be used to distinguish them from GR (see Pani and Cardoso (2009); Cardoso and Gualtieri (2009); Molina et al. (2010); Pani et al. (2013a, b); Blázquez-Salcedo et al. (2016, 2020a, 2020b); Pierini and Gualtieri (2021); Wagle et al. (2022); Cano et al. (2022); Pierini and Gualtieri (2022); Cano et al. (2023); Wagle et al. (2024); Antoniou et al. (2024); Antoniou (2024); Cano et al. (2024); Chung and Yunes (2024a, b) for recent progress in computing QNMs in theories beyond GR). In particular, any modified gravity theory generically introduces two modifications to the ringdown: i) a deformed QNM spectrum, whose deviations from the GR case are proportional to (powers of) the coupling constant; ii) novel extra modes in the gravitational waveform D’Addario et al. (2024); Crescimbeni et al. (2024), whose amplitude is proportional to (powers of) the coupling.
@@
BHs for $p>0.876$, instead, are stable for radial perturbations. Time-domain evolutions of BHs in the Schwarzschild branch Held and Lim (2021), give strong indications for their non-radial stability (see also Held and Lim (2023) for the evolution of rotating BHs and of BH binaries). Moreover, scalar and electromagnetic perturbations of BHs in the non-GR branch have been studied using approximate techniques, finding stable modes Cai et al. (2016); Zinhailo (2018, 2019); Konoplya (2023). In this paper we compute for the first time gravitational QNMs, in the axial sector, for both branches of spherically symmetric BH solutions in EW gravity, finding as expected a deformation of the GR spectrum and new classes of modes.
Moreover, we study whether the new modes can be excited in the ringdown stage of a binary BH coalescence. Then, using the results of our study, we discuss the phenomenological perspectives of EW gravity, i.e. whether we expect the non-GR features of BHs in this theory, if present, could be observed by present or near-future GW detectors.
The paper’s structure is the following: in Sec. II we present the theoretical framework of the model. In Sec. III we review the background solutions in this theory, using both numerical and semi-analytical treatments. In Sec. IV we analyze the axial perturbations in EW gravity, first on a Schwarzschild background recovering the results presented in Brito et al. (2013). We show our analysis for the axial QNMs of the non-GR branch of solutions in Sec. IV.2. In Sec. V we discuss the phenomenological perspectives of EW gravity. Finally, we draw our conclusions in Sec. VI.
@@ II Framework
where the GB invariant is defined as $\mathscr{G}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. The dimensionful parameters $\alpha$ and $\beta$ are associated with additional massive degrees of freedom. Specifically, in addition to the massless spin-2 graviton, this theory predicts a massive spin-0 mode with mass $m^{2}=1/(6\beta)$, and a massive spin-2 mode with mass $\mu^{2}=1/(2\alpha)$. Remarkably, the sign of the kinetic term of the massive spin-2 mode is the opposite than those of the other kinetic terms, i.e. it is a ghost-like degree of freedom. This suggests that quadratic gravity is non-unitary as a full quantum theory (however, see Donoghue and Menezes (2022) for different interpretations); this problem is avoided by considering quadratic gravity as a low-energy EFT, as long as the energy scale is lower than a cut-off scale which is of the order of $\alpha^{-1/2}\sim\mu$. However, in this framework the actual deviations from GR are expected to occur at higher order in the EFT expansion, since the quadratic corrections can be transformed by a field redefinition to higher-order terms Burgess (2004); Endlich et al. (2017).
Note that the existence of the hairy BH solution, occurring for $p=r_{h}\mu\sim 1$ and then at energy scales close to $\mu$, is questionable in an EFT framework. For completeness, in this work we will compute the QNMs of both the Schwarzschild BH solution (which clearly exists also in the EFT) and of the hairy BH solution (which instead exists only in the full theory). We also remark that the existence of an Ostrogradsky ghost does not prevent the theory from having a well-posed initial value formulation at a classical level Noakes (1983); Held and Lim (2021).
@@
where beyond-GR contributions are proportional to a factor $\alpha$. Despite not being obvious at first glance, the field equations for spherically symmetric configurations yield second-order equations. We will go over that in the following subsections.
@@ II.1 Auxiliary field
$\begin{split}&\mathcal{E}^{(f)}_{\mu\nu}\equiv G_{\mu\nu}+G_{(\mu}{}^{\rho}f_{\nu)\rho}-g_{\mu\nu}G^{\rho\sigma}f_{\rho\sigma}-fR_{\mu\nu}\\ &+Rf_{\mu\nu}+\Box f_{\mu\nu}+(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu})f-2\nabla_{\rho}\nabla_{(\mu}f_{\nu)}{}^{\rho}\\ &+g_{\mu\nu}\nabla_{\rho}\nabla_{\sigma}f^{\rho\sigma}+\mu^{2}\big{[}\big{(}f^{2}-f_{\rho\sigma}f^{\rho\sigma}\big{)}/2\\ &+2\big{(}f_{\mu}{}^{\rho}f_{\nu\rho}-f\,f_{\mu\nu}\big{)}\big{]}=0\,.\end{split}$ (10)
We remark that the GR limit corresponds to $\mu\to\infty$, i.e. $\alpha\to 0$. In this limit, the auxiliary field $f_{\mu\nu}$ (Eq. (8)) vanishes, and Eqs. (10) reduce to Einstein’s equations in vacuum. Conversely, in the $\mu\to 0$ limit the theory becomes strongly coupled, the action only has quadratic curvature terms, and EW gravity reduces to the pure Weyl gravity Weyl (1918).
@@ III Background solutions
From Eq.,(9), using the symmetry of $f_{\mu\nu}$, we find $\nabla_{\mu}\nabla_{\nu}f^{\mu\nu}=\nabla_{\mu}\nabla_{\nu}f^{\nu\mu}=\Box f$. From the trace of (10) we find then $R=0$: the spherically symmetric solutions (in vacuum) of EW gravity are necessarily Ricci-scalar flat. This also means that from (7) the trace of the auxiliary field vanishes, i.e. $f=0$. From this property we derive the first equation for the metric functions, i.e.
@@
Given that the static and spherically symmetric solutions have necessarily $R=0=f$, we will divide the analysis of BH solutions in the following two cases:
@@
Obviously in the first case the solution can be written in analytical form, while the in the second case it has to be found numerically, by solving the system of ordinary differential equations with suitable boundary conditions.
@@ III.1 Expansions at the horizon
To find the coefficients in the expansions we substitute them in the background equations, which we then solve order by order. We show here the first few coefficients:
@@ III.3 Semi-analytical background
In Fig. 1 we present, as an example, the $(tt)$ and $(rr)$ metric components as functions of $r/r_{h}$, for the hairy BH solution with $p=1.1$. We show both the solution computed solving numerically the field equations, and the semi-analytical approximation up to fourth order in terms of continued fractions. For the numerical solution the process we follow is a shooting approach with respect to the non-GR parameter $b_{1}$, integrating from the horizon outwards, so that at asymptotic infinity the exponentially diverging solution vanishes. For comparison, we also show the Schwarzschild solution.
@@
III.4 Domain of existence
The branch of non-GR BH solutions can be found for Kokkotas et al. (2017)
@@ III.4 Domain of existence
where we remind the reader that the dimensionless parameter $p$ is defined as $p\equiv{r_{h}}/{\sqrt{2\alpha}}\equiv r_{h}\mu$ In Fig. 2 we show the domain of existence and of radial stability of hairy BHs in EW gravity, corresponding to the decreasing curve. The vertical dashed line highlights the allowed range, which starts at the bifurcation point from the Schwarzschild branch, and ends at the point where solutions have vanishing mass. The straight line, instead, corresponds to the solutions of the Schwarzschild branch.
@@
Figure 2: Domain of existence of BHs in EW gravity. The bifurcation point from the GR solution is found at $p_{\text{min}}\simeq 0.876$, while the the upper value of the positive mass range is located at $p_{\text{max}}\simeq 1.143$. The dashed part of the curves corresponds to solutions either radially unstable, or having negative mass.
@@ IV Perturbations
where $\varepsilon$ is a bookkeeping parameter, $\bar{g}_{\mu\nu}$ and $\bar{f}_{\mu\nu}$ are the background metric and auxiliary field corresponding to a static, spherically symmetric BH solution, and $\delta g_{\mu\nu}$, $\delta f_{\mu\nu}$ are their respective perturbations. Hereafter, any quantity evaluated on the non-perturbed background will be denoted with a bar. At first order in the perturbations, Eq. (6) gives:
@@ IV.1 Schwarzschild background
From the above, it is evident that the massive perturbations $\delta f_{\mu\nu}$ decouple from $\delta g_{\mu\nu}$. Indeed, setting $\delta f_{\mu\nu}=0$, Eq. (34) is satisfied, and Eq. (33) reduces to the standard Regge-Wheeler equation Regge and Wheeler (1957), leading to the QNMs of Schwarzschild BHs in GR. When $\delta f_{\mu\nu}\neq 0$, Eqs. (33)-(34) yield a set of equations for the massive spin-two degrees of freedom, which coincide with those presented in Brito et al. (2013); Ould El Hadj (2024) after substituting the Schwarzschild metric. Although those perturbations are massive and do propagate at infinity, they nevertheless source ordinary metric perturbations through Eq. (33). Axial perturbations of the massive spin-2 excitation $\delta f_{\mu\nu}$ can be decomposed in spherical tensor harmonics (see Appendix B), leading - for each value of the harmonic indexes - to two vector perturbations, $F_{0}(r)$, $F_{1}(r)$, and one tensor perturbation, $F_{2}(r)$. In the massless limit, the only dynamical perturbation in the decomposition of $\delta f_{\mu\nu}$ is the tensor perturbation $F_{2}$, while the vector ones become pure gauge  [^1] Note that in this limit one does not recover, for $\delta f_{\mu\nu}$, the same decomposition as the (axial) massless field $\delta g_{\mu\nu}$, because the tensorial decomposition of $\delta f_{\mu\nu}$ is different from the Regge-Wheeler gauge employed in the decomposition of $\delta g_{\mu\nu}$. In particular, on a Schwarzschild background $\delta f_{\mu\nu}$ is traceless and transverse. .
@@
$\big{[}r\left(r\omega^{2}-BA^{\prime}\right)-A\left(rB^{\prime}-2B+\mu^{2}r^{2}+\Lambda-2\right)\big{]}$
$=0\,,$ (38)
where $\Lambda\equiv\ell(\ell+1)$. Finally, from Eq. (33), by eliminating $h_{0}$, we find
@@
We may now eliminate $F_{0}$ by making use of (35), in which case the decoupled equations for the massive perturbations can be written as a system of Schrödinger-type equations
@@
The terms $-A\mu^{2}$ in $\hat{V}_{11}$ and $\hat{V}_{22}$ are an indication that $Q$ and $Z$ correspond to massive modes propagating with mass $\mu^{2}$. Notice that with these redefinitions for $F_{1}$ and $F_{2}$ we cannot simultaneously describe all three perturbation functions $(F_{1},F_{2},h_{1})$ in a Schrödinger-type system due to the presence of the $F_{2}^{\prime}$ term in (39).
In order to study the full QNM spectrum, we solved the following system describing both massless and massive degrees of freedom:
@@
For ingoing waves at the horizon, ${\Psi}^{(j)}(r)\sim e^{-i\omega r_{*}}$. Therefore, we expand
$h_{1}(r)=$ $\;e^{-i\omega r}(r-2M)^{-2iM\omega}\sum_{n=0}\tilde{h}_{1}^{(n)}(r-2M)^{n-1}\,,$ (47)
@@
Summarizing, for each value of the BH mass (i.e. of the horizon radius), for each $\ell$, we find three series of modes with axial parity, parametrized as customarily by the principal number $n$: (i) the massless modes, which coincide with those of GR; (ii) the massive vector modes; (ii) the massive tensor modes.
We present our results in Figs. 3 4, where we show the fundamental ($n=0$) $\ell=2$ massive vector and tensor modes, respectively, computed using the DI approach with $n_{\text{max}}=4,8$, and the CF approach. We find that the DI results, increasing $n_{max}$, converge towards the CF results, which we use as a benchmark. The vector mode computed using the DI approach with $n_{\text{max}}=4$ is a very good approximation, within $1\%$, of the CF result for $p\lesssim 0.6$, while for larger values of $p$ we need $n_{\max}=8$ to reach the same level of accuracy. For the tensor mode, the DI approach with $n_{\text{max}}=4$ is accurate within $1\%$ for $p\lesssim 0.6$, while that with $n_{\text{max}}=8$ has the same accuracy for $p\lesssim 0.75$. For larger $p$ the error increases but is smaller than $5\%$ for $p\lesssim 0.95$. In Fig. 4 we also show the fundamental massless mode, which also belongs to the spectrum, and is the same as in GR. It coincides with the $\mu\to 0$ limit of the massive tensor mode, while we remind that the massive vector mode is not dynamical in this limit (see Appendix B).